Rita rows 12 km upstream and 12 km downstream in 3 hours. The speed of her boat in still water is 9 km/hr. Find the speed of the stream
Answer by checkley71(8403)

12/(9-X)+12/(9+X)=3
[12(9+X)+12(9-X)]/(9+X)(9-X)=3
[108(12X+108-12X]/(81-X^2)=3
216/(81-X^2)=3 NOW CROSS MULTIPLY
3(81-X^2)=216
243-3X^2-216=0
-3X^2+27=0
-3(X^2-9)=0
-3(X+3)(X-3)=0
X-3=0
X=3 KMH IS THE SPEED OF THE CURRENT.
PROOF:
12/(9-3)+12/(9+3)=3
12/6+12/12=3
2+1=3
3=3